log[6](x+1)+log[6](x)=1
Answer by lwsshak3(11628)

Solve for x
log[6](x+1)+log[6](x)=1
log6[x(x+1)]=1
convert to exponential form: base(6) raised to log of the number(1)=number[x(x+1)]
6^1=[x(x+1)]
6=x^2+x
x^2+x-6=0
(x+3)(x-2)=0
x=-3(reject, (x+1)>0)
or
x=2